Two boats left an island at the same time to return to their home port. Their average speeds differ by 4MPH. The faster boat reaches the port after 3 hours. The slower boat took thirty minutes longer. Find the speed of each boat?
Answer by stanbon(75887)

Two boats left an island at the same time to return to their home port.
Their average speeds differ by 4MPH. The faster boat reaches the port after 3 hours. The slower boat took thirty minutes longer. Find the speed of each boat?
--------------
Faster boat DATA:
time = 3 hrs ; rate = x mph ; distance = rt = 3x miles
--------------------------
Slower boat DATA:
time = 3.5 hrs ; rate = (x-4)mph ; distance = rt = 3.5(x-4) miles
--------------------------
Equation:
distance = distance
3x = 3.5x-14
-0.5x = -14
x = 28 mph (speed of the faster boat)
x-4 = 24 mph (speed of the slower boat)
==========================================
Cheers,
Stan H.